OpenStudy (jiteshmeghwal9):

If \(\sqrt{\dfrac{x}{1-x}}+\sqrt{\dfrac{1-x}{x}}=2\dfrac{1}{6}\) the values of \(x\) are :-

OpenStudy (experimentx):

let one be 'a' the other one is '1/a' ... then solve for 'a'

OpenStudy (jiteshmeghwal9):

I did this as follows :- \[(\sqrt{x/1-x})^2+(\sqrt{1-x/x})^2+2(\sqrt{x/-x})(\sqrt{1-x/x})=(13/6)^2\]

OpenStudy (hba):

Substitution is the key my friend.

OpenStudy (anonymous):

not as bad as it looks i think i added on the left

OpenStudy (jiteshmeghwal9):

@experimentX will u give me the equation & @satellite73 how did u add on left ?

OpenStudy (anonymous):

\[\frac{1}{\sqrt{x(1-x)}}=\frac{13}{6}\]

OpenStudy (jiteshmeghwal9):

@hba how can we substitute it ?

OpenStudy (experimentx):

\[ a + \frac 1 a = \frac{13} 6\]

OpenStudy (jiteshmeghwal9):

LEMME TRY

OpenStudy (anonymous):

then \[\sqrt{x(1-x)}=\frac{6}{13}\] etc

OpenStudy (jiteshmeghwal9):

X-X^2=6/13 LEFT IS COMPLETING SQUARE, RIGHT ?

OpenStudy (jiteshmeghwal9):

36/169*

OpenStudy (experimentx):

also you can take LCM, i guess you 'll get the result same as satellite did.

OpenStudy (jiteshmeghwal9):

ohh ! i gt it now thanx :)

OpenStudy (anonymous):

i bet there is a snazzier way, i just grind it till i find it, but i would like to see snapper solution